OpenStudy (anonymous):

Simplify this:

OpenStudy (anonymous):

\[\frac{ x ^{2}-9 }{ 4x-12 }-\frac{ x+3 }{ 2 }\]

OpenStudy (john_es):

\[\frac{(x+3)(x-3)}{4(x-3)}-\frac{x+3}{12}=\frac{x+3}{4}-\frac{x+3}{12}=\] \[=\frac{(x+3)}{4}(1-1/3)=\frac{x+3}{6}\]

OpenStudy (anonymous):

the answer i got here is 1/2

OpenStudy (cwrw238):

john's solution is correct the answer is only 1/2 if x = 0

OpenStudy (anonymous):

ok thanks